CH/DOC/93/02

Report

Legnaro Meeting

Author: E. Plagnol

Date: 5 Feb 1993

Summary of the Initial Meeting of a Forum on Nuclear Reaction Mechanisms Legnaro, 16-18 December 1992

As a result of informal invitations, a group of physicists involved in the general area of nuclear reaction mechanism studies and whose interest range from energies close to the Coulomb barrier to several GeV/N met in Legnaro. The meeting examined ways by which closer links between groups working in this field could be enhanced for the general benefit.

It became clear during the two days of discussions that there was a vast reservoir of information and expertise from which the whole community could benefit if it could be disseminated effectively. The information ranges from detection techniques used through nuclear models required to interpret the results.

The meeting identified a pressing need to:

• Identify the sources of expertise,

• Examine ways to effectively exchange the information and knowledge,

• Encourage closer cooperation between members of our community.

To satisfy these needs, the meeting decided to try out two initiatives during 1993:

• A compilation of “information sheets” based on existing and future programs. These would describe, among other things detector systems, associated expertise, available simulation computer codes and event generators.

• The encouragement of small informal gathering of experts to work on specialized topics. These would bring together a limited (10) number of people in an environment with such technical support that would allow them to work efficiently.

More information on these initiative for 1993 can be found at the end of this summary. It was recognized by those present that many groups working in this field were not represented either because they were unable to attend or because of the informal way the meeting was convened.

The meeting emphasised that the forum is open to all. Individuals, groups or institution that wish to be informed, to take part actively, suggest other subjects of discussions or present other forms of cooperation are welcomed and should reach any of the following “contact persons” : F.Gramegna (Italy), A.Gobbi (Germany, to be confirmed), D.L.Watson (UK), Z.Majka (Poland) and E.Plagnol (France ; coordinator).

The meeting has stated very clearly that the forum’s goals and activities must fit in perfectly with those of existing bodies or institutions such as NUPECC, the EPS, and the local physical societies.

Contact between the forum and these institutions will be established shortly. In order to appraise the response of the community and to discuss in more depth

and details the long term development of the forum, two meetings will be held at the middle and at the end of 1993.

The forum’s activity in 1993

Two main initiatives have been suggested for 1993 :

• The first one consists in the compilation of ”information sheets”. Two types of information sheets are being considered for the moment :

Existing and future projects or collaborations. The content should insist on the description of the detectors that have been or are developed and on the experts (name, telephone,fax,...) that contributed to it. The persons responsible for this are JP.Coffin (CRN - Strasbourg) and U.Lynen (GSI - Darmstadt).

A description of existing event generators computer codes and detector simulation programs. The persons responsible for this are G.Casini (INFN - Florence) and D.Guinet (IPN - Lyon).

Z.Majka (Jag. Univ. - Krakow) is responsible for the collect of information on these subjects in Eastern Europe.

• The second is the organisation of “working group meetings”. These meetings are intended to group a limited number of specialists (of the order of 10), on a given well defined topic, for a limited period of time (of the order of one week) in a location having the relevant technical support (computer, electronics, library,…) such that they can work efficiently on the subject and exchange information and knowledge at a technical level. In 1993 5 subjects have been selected :

– The performances of Silicon-Cesium Iodide telescopes (resp. : A.Moroni and D.L.Watson),

– Timing with large area Silicon detectors (resp. : F.Porto ), – A study of existing event generators around the Fermi energy (E.Plagnol

and Z.Majka),

– A comparison between various event generator computer codes (M.Baldo),

– The status of collaboration with Eastern European groups (A.Gobbi).

As can be seen, the subjects cover a wide range of topics and concern different levels of activity. People who are interested in discussing these subjects and the way they can be best treated are welcomed to contact the people in charge.

Parallel to these activities, G.Auger will look into the possibility of establishing a computer based “news centre” where the relevant informations concerning the forum could be concentrated and looked up by members of the community. This could eventually evolve into a data support for the “information sheets” mentioned above.

Postscript

Dear Friend, After some delay, we are able to send you the summary of our meeting in Leg-

naro.

If you were not present at this meeting, we hope that you will help us to give this initiative some “momentum”. The best way to do this is

(i) to spread the information around you, (ii) to talk to the directors of your lab to ask how you could cooperate, and (iii) mainly suggest news ideas to make this initiative a success. If you would like to ogranize a meeting yourself or if you want to participate in

one of the meetings already planned, please contact the organisers.

Thank you for your cooperation. Eric Plagnol

Figure 1: List of participants :

Name  Institution  E-Mail
M.BRUNO  UNIV. OF BOLOGNA BRUNO@BO.INFN.IT
A.MORONI  INFN - MILANO  AMORONI@MILANO.INFN.IT
D.L.WATSON  UNIV. OF YORK  OE@PHGA.YORK.AC.UK
J.S.LILLEY  DARESBURY LAB.  BJ@NNGA.DARESBURY.AC.UK
F.GRAMEGNA  L.N.L - LEGNARO  FABY@LNL.INFN.IT
Z.MAJKA  JAG. UNIV. KRAKOW UFMAJKA@PLKRCY11
E.MIGNECO  LNS - CATANIA  MIGNECO@LNS.INFN.IT
U.LYNEN  GSI - DARMSTADT  LYNEN@DDAGSI5
F.SAINT-LAURENT GANIL - CAEN  STLAUREN@FRCPN11.IN2P3.FR
F.RAMI  CRN - STRASBOURG RAMI@FRCPN11.IN2P3.FR
A.GOBBI GSI - DARMSTADT GOBBI@DDAGSI5
JP.COFFIN  CRN - STRASBOURG COFFIN@FRCPN11.IN2P3.FR
G.CASINI  INFN - FLORENCE  CASINI@FI.INFN.IT
M.SANDOLI  UNIV. NAPOLI  SANDOLI@NA.INFN.IT
CH.LE BRUN  LPC - CAEN  LEBRUNCH@FRCPN11.IN2P3.FR
P.SAPIENZA  LNS - CATANIA  PIERA@.LNS.INFN.IT
A.FERRERO  INFN - MILANO  FERRERO@MILANO.INFN.IT
M.D’AGOSTINO  UNIV. OF BOLOGNA DAGOSTINO@BO.INFN.IT
M.STERN  IPN - LYON  STERN@FRCPN11.IN2P3.FR
D.GUINET  IPN - LYON  GUINET@FRCPN11.IN2P3.FR
F.PORTO  LNS - CATANIA  PORTO@VAXLNS.LNS.INFN.IT
M.BALDO  INFN - CATANIA  BALDO@CATANIA.INFN.IT
G.VANNINI  UNIV. TRIESTE  VANNINI@TS.INFN.IT
N.GELLI  UNIV. OF FIRENZE  GELLI@FI.INFN.IT
M.CINAUSERO  INFN - FIRENZE  CINAUSERO@FI.INFN.IT
J.C.STECKMEYER LPC - CAEN  STECK@FRCPN11.IN2P3.FR
G.AUGER  GANIL - CAEN  AUGER@FRGAN01.IN2P3.FR
E.PLAGNOL  GANIL - CAEN  PLAGNOL@GANAC4.IN2P3.FR
R.A. RICCI  L.N.L. -LEGNARO
G. MARGAGLIOTTI UNIV. TRIESTE  MARGAGLIOTTI@TRIESTE.INFN.IT
L. VANNUCCI L.N.L. LEGNARO VANNUCCI@LNL.INFN.IT
U. ABBONDANNO INFN TRIESTE  ABBONDANNO@TS.INFN.IT
R. MORO  UNIV. NAPOLI  MORO@NA.INFN.IT
N. COLONNA  INFN BARI  COLONNA@BA.INFN.IT
P.M. MILAZZO  UNIV. BOLOGNA  MILAZZO@BO.INFN.IT
L. FIANDRI  UNIV. BOLOGNA  FIANDRI@BO.INFN.IT

Figure 2: The following people have expressed their interest in the forum but were not able to come this time:

I. IORI  UNIV. MILANO  IORI@MILANO.INFN.IT
V. PATICCHIO  INFN BARI  PATICCHIO@BA.INFN.IT
A. PANTALEO  INFN BARI  PANTALEO@BA.INFN.IT
E. FIORE  INFN BARI  FIORE@BA.INFN.IT
F. LUCARELLI  UNIV. FIRENZE  LUCARELLI@FI.INFN.IT
A. STEFANINI  UNIV. FIRENZE  STEFANINI@FI.INFN.IT
P. MAURENZIG  UNIV. FIRENZE  MAURENZIG@FI.INFN.IT
G. LARANA  UNIV. NAPOLI  LARANA@NA.INFN.IT
E. FIORETTO  UNIV. NAPOLI  FIORETTO@NA.INFN.IT
B.R. FULTON  UNIV. BIRMINGHAM BW@NP.PH.BHAM.AC.UK
T. REPOSEUR  NANTES  REPOSEUR@FRCPN11.IN2P3.FR
D. ARDOUIN  NANTES  ARDOUIN@FRCPN11.IN2P3.FR
C. BECK  STRASBOURG  BECK@FRCPN11.IN2P3.FR
F.P. HESSBERGER GSI - DARMSTADT KC50@DDAGSI3
NINOV  GSI - DARMSTADT NASE@DDAGSI5
J. GERL  GSI - DARMSTADT GERL@DDAGSI5
MORGERNSTERN BERLIN  MORGERNSTERN@VAX.HMI.DBP.DE
R. SIEMSSEN  GRONINGEN  SIEMSSEN@GANAC4.IN2P3.FR
FERRERO  IFIC VALENCIA  FAX: 346-3864583
P. DIAZ  IFIC VALENCIA  16444::DIAZ
J.A. RUIZ  IFIC VALENCIA  16444::JAR